# Learning on the Fly: Rapid Policy Adaptation via Differentiable Simulation

### v1

@@ Abstract
Learning control policies in simulation enables rapid, safe, and cost-effective development of advanced robotic capabilities. However, transferring these policies to the real world remains difficult due to the sim-to-real gap, where unmodeled dynamics and environmental disturbances can degrade policy performance. Existing approaches, such as domain randomization and Real2Sim2Real pipelines, can improve policy robustness, but either struggle under out-of-distribution conditions or require costly offline retraining. In this work, we approach these problems from a different perspective. Instead of relying on diverse training conditions before deployment, we focus on rapidly adapting the learned policy in the real world in an online fashion. To achieve this, we propose a novel online adaptive learning framework that unifies residual dynamics learning with real-time policy adaptation inside a differentiable simulation. Starting from a simple dynamics model, our framework refines the model continuously with real-world data to capture unmodeled effects and disturbances such as payload changes and wind. The refined dynamics model is embedded in a differentiable simulation framework, enabling gradient backpropagation through the dynamics and thus rapid, sample-efficient policy updates beyond the reach of classical RL methods like PPO. All components of our system are designed for rapid adaptation, enabling the policy to adjust to unseen disturbances within 5 seconds of training. We validate the approach on agile quadrotor control under various disturbances in both simulation and the real world. Our framework reduces hovering error by up to 81% compared to -MPC and 55% compared to DATT, while also demonstrating robustness in vision-based control without explicit state estimation.
Index Terms:
Machine Learning for Robot Control, Aerial Systems: Perception and Autonomy, Continual LearningI Introduction
@@
Robot learning through simulation has seen great success in recent years, thanks to the rapid improvements in computer hardware and advancements in efficient physics simulation [1]. Simulation provides a fast, safe, and cost-effective way to collect data and train policies, enabling experiments that would be impractical or unsafe in the real world. However, transferring control policies learned purely in simulation to physical systems is challenging. While a high-fidelity simulation model may be used, the system parameters are often difficult to precisely identify. In addition, unmodeled effects such as aerodynamic turbulence, sensor noise, and actuator delays further complicate the real-world dynamics, thus making accurate alignment between simulation and reality difficult to achieve. The resulting mismatch, known as the sim-to-real gap [2], remains a central obstacle to deploying learning-based controllers in the real world. Bridging this gap is essential to retain the advantages of simulation while ensuring that policies perform reliably under the complexity and variability of real-world conditions.
@@
Domain randomization is a common strategy to address this issue [3], in which simulation parameters such as dynamics [4, 5], sensor noise [6, 7], and visual appearance [3, 8] of the environment are varied during training to expose the policy to a wide range of possible deployment scenarios. By learning in diverse conditions, the agent develops robust policies that are less likely to overfit to the specific characteristics of a single environment. However, domain randomization cannot exhaustively anticipate all possible real-world conditions. Thus, when the environment shifts beyond the randomized distribution, policy performance will strongly degrade to out-of-distribution disturbances [9]. Beyond domain randomization, Real2Sim2Real methods [6, 5] have shown strong sim-to-real transfer capabilities through offline refinement of the simulation model using real-world data before retraining policies for deployment. While such methods are effective in improving transfer, they require extensive data and costly retraining. For example, [10] reports 75 minutes of real-world data collection, which makes them unsuitable for rapid online adaptation to changing conditions.
@@
In this work, we approach these problems from another perspective: we propose to rapidly adapt the policy to unknown external disturbances in the real world, in an online fashion. The core insight of the proposed framework is to integrate online residual dynamics learning with rapid policy adaptation via differentiable simulation. All system components are designed to update both the residual dynamics and policy as quickly as possible, ideally within a few seconds, during runtime. In this way, the policy becomes adaptive by continuously “overfitting” rapidly to the current environment scenario, and paradoxically, more “generalizable” across diverse real-world conditions.
For our pipeline, we start with a lightweight rigid-body dynamics model and continuously refine it by learning residual dynamics from real-world flight data. The residual-augmented dynamics model is embedded in a differentiable simulation framework to achieve more accurate and sample-efficient policy adaptation. Differentiable simulation provides the key advantage here: by allowing accurate gradients to flow through the dynamics, it makes real-world policy adaptation more efficient than classical RL approaches such as PPO [11].
@@
Another key innovation is our alternating optimization scheme, where policy learning and residual model learning are interleaved so that each batch of real-world data is used efficiently for both dynamics refinement and control improvement through simulation using the refined dynamics. All of these components ensure the simulation is aligned with reality and enables rapid, efficient policy adaptation to unknown disturbances, even controlling directly from perceptual input.
@@
We evaluate our proposed framework in both simulation and real-world experiments, considering a range of environmental disturbance conditions on an agile quadrotor platform, whose nonlinear dynamics and sensitivity to aerodynamic effects make it a challenging benchmark for adaptive control [12]. In state-based control tasks such as hovering, where the policy receives as input full quadrotor state information, our method attains an average error of , an reduction over -MPC ( ) and over DATT [13] ( ), while ensuring stable flight under modeling errors and out-of-distribution disturbances. In visual feature-based control, our framework achieves similar gains, demonstrating that rapid adaptation remains effective under partial or noisy observations - a capability unattainable with classical control methods in the absence of state estimation.
@@
We propose an online policy adaptation framework that combines residual dynamics learning with differentiable simulation, enabling real-world policy adaptation to unknown disturbance within 5 seconds of training. We propose an alternating optimization scheme that interleaves dynamics learning and policy learning, ensuring that real-world samples effectively improve both components. The framework supports both state-based and visual feature-based inputs (without explicit state estimation), and we demonstrate its effectiveness through simulation experiments and real-world validation, where it outperforms both classical and learning-based controllers under large unseen disturbances. Together, our framework shows that policies can learn and adapt within seconds in the real world, reducing the reliance on domain randomization, which can fail to capture real-world complexity and adapt to out-of-distribution scenarios.
@@ II-A Aligning Simulation with the Real-World
Closing the sim-to-real gap requires quantifying the misalignment between simulation and real-world dynamics, typically through system identification or residual dynamics learning. System identification estimates parameters of an analytical dynamics model from input–output data [14], but its representation capacity is limited to the modeled system [15], making it less effective for capturing complex dynamics and disturbances. Residual dynamics learning addresses this by directly modeling the discrepancy between analytical predictions and real-world measurements. It has been applied to improve quadrotor odometry and tracking [10, 16], learn motor delays in quadrupeds [6], and predict residual forces in soft robots [17].
@@ II-B Fast Policy Learning in Simulation
While traditional RL methods suffer from prohibitively long training times, significant speed-ups can be achieved using highly optimized physics simulation [18]. However, these methods are limited by sample inefficiency due to the high variance in the zeroth-order policy gradient estimates. A recent alternative paradigm, policy learning via differentiable simulation, uses smooth, differentiable dynamics and rewards to enable policy learning via first-order gradients [19], offering substantial gains in sample efficiency and training time over RL [20]. It has been applied to direct policy parameterizations, such as parametric curve frequencies for swimming robots [21] and sinusoidal policies for robotic cutting [22]. For applications to neural network policies, however, unstable gradients often limit applications to short-horizon tasks with simplified contacts and restricted start-state variation [23, 24]. To address this, prior work has explored enhancements such as early-stopping simulations at contact events, truncated BPTT [25], and reward augmentation with a learned critic [26, 27].
@@ II-C Learning-Based Adaptive Control
Residual dynamics models have been used for online disturbance estimation, via offline-trained networks [28], Gaussian Processes [16], or differentiable simulation–based system identification [29]. However, these methods mainly augment optimization-based controllers like MPC, which rely on full state information, and do not directly extend to vision-based control. Neural network policies have also been conditioned on disturbance estimates [13, 30, 4], but since they are trained offline in randomized simulations and remain fixed during deployment, they struggle with domain shifts and unseen real-world conditions [31, 13].
@@ III Methodology
Our approach consists of two phases: policy pretraining and online adaptation. During pretraining, we train a base policy using a simplified, low-fidelity analytical dynamics model without residual dynamics, which is used as the initial policy for online adaptation. During online adaptation (see Fig. 3), residual dynamics learning, policy adaptation, and real-world deployment run in parallel across multiple threads, with parameters exchanged efficiently between processes via ROS as serialized byte strings. The residual dynamics network is continuously updated from a rolling buffer of quadrotor states and actions, and combined with the analytical model to form a hybrid dynamics embedded in the differentiable simulation for policy adaptation. Finally, the deployment loop uses the latest policy network parameters to continuously output control commands at 50 to the on-board flight controller, and simultaneously collects flight trajectories.
@@ III-A Differentiable Simulation Model
We model the quadrotor as a discrete-time dynamical system with continuous state and action spaces and , respectively. The system evolves according to the differentiable hybrid dynamics model which comprises the analytical and learned residual components, and describes the system evolution over time. At time step , an observation model generates an observation from the state , and is passed as input to a deterministic and differentiable policy network which outputs an action , and finally a deterministic, smooth and differentiable reward function emits a reward based on the state-action pair. Thus, all components are fully-differentiable and allows gradient backpropagation through the simulation.
@@ III-C Policy Optimization Using Analytical Gradients
By leveraging the differentiable dynamics and reward structure, we can obtain first-order analytical policy gradients of the objective (2) via Back-Propagation Through Time (BPTT) (see [32] for a full derivation). The gradient and the update rule of the policy parameters are given by
where is the derivative matrix of the system dynamics , and is the learning rate. We build upon an existing open-source differentiable simulator for quadrotors [20] written entirely in JAX to leverage both its automatic-differentiation framework for computing the analytical policy gradients and GPU-accelerated computing for efficient parallel simulation.
@@ III-D Residual Dynamics Learning
Given the concatenated input vector of quadrotor state and action , an MLP network parameterized by is trained to predict the residual acceleration , defined as the difference between the ground-truth acceleration measured on the real system and the theoretical acceleration from the analytical dynamics in (1). The residual acceleration training targets are computed as . Given a batch of samples , we train the model by minimizing the loss function via where is the weight matrix of the -th network layer, and controls the regularization strength. The loss comprises a standard MSE term and a spectral norm regularization term, where the latter has been shown to improve generalization beyond the training distribution [33] by regulating the network’s Lipschitz constant [28]. We minimize the loss using the Adam optimizer [34].
@@ III-E Design Choices for Maximum Runtime Efficiency
During forward simulation, we use a hybrid dynamics model obtained by additively combining the analytical and learned residual dynamics models. Here, we use a simple, low-fidelity analytical dynamics model (see Sec. III-B) which models the quadrotor as a point-mass. The resulting acceleration given a state and action input pair is computed as , where is the network prediction. The quadrotor states are simulated at 50 via Runge-Kutta 4 time-integration of the dynamics using . While the hybrid dynamics model composed of differentiable analytical and learned components remains overall fully differentiable, we only perform gradient backpropagation through the analytical dynamics model and not the frozen network to obtain the policy gradients. This was inspired by prior work in policy learning using differentiable simulation for both quadruped [11] and quadrotor [20] control, which showed that combining accurate forward dynamics simulation with the backpropagation of a surrogate gradient based on a simplified dynamics model achieves faster runtime without impacting the resulting policy performance. We analyze and justify the above design choices through simulated experiments, and present and discuss the results in Sec. IV-C.
@@
We compare two existing methods of adapting a pretrained policy: full vs. low-rank adaptation (LoRA) [35]. Full adaptation involves updating all parameters of the policy network, similar to [36], whereas LoRA freezes all pretrained parameters and instead adapts an additive low-rank network module [37] which forms a lower-dimensional trainable parameter space. The latter has been shown to achieve more memory and parameter-efficient policy adaptation using RL for task-transfer [35] and multi-agent [38] learning. Therefore, as an exploratory comparison, we seek to understand whether LoRA can also be combined with sample-efficient policy learning using differentiable simulation to effectively adapt a pretrained policy to unknown environmental disturbances.
@@ IV-A1 Task and Reward Definitions
We evaluate our approach on stabilizing hover and trajectory tracking tasks for the quadrotor platform. For stabilizing hover, the policy is required to regulate the quadrotor state towards a goal position and maintain it at all times, which is non-trivial given the quadrotor’s non-linear and unstable dynamics. We evaluate both a state-based control policy which at each time step receives the observation , and an end-to-end visual feature-based policy which only receives the projected pixel coordinates of seven 3D keypoints from the past five time steps and the last three actions. For real-world experiments, the 3D keypoints are simulated in a hardware-in-the-loop style using quadrotor state estimates from a motion-capture system. Our training setup closely follows the open-source environment setup in [20]. Trajectory tracking instead requires following a reference trajectory defined as a time-parameterized sequence of quadrotor states, which is generated to be smooth up to the acceleration level. We use two such trajectories, Circle and Figure-8, for this task in a state-based training setting (see Fig. 4). For both tasks, the reward at each time step is defined as a sum of position, velocity, and actuation reward terms For stabilizing hover, the individual reward terms are given by , and where is the Huber loss, and is the mass-normalized action required to counteract gravity. For trajectory tracking, the individual reward terms are given by and where and are respectively the corresponding position and velocity of the reference trajectory at time .

### v2

@@ Abstract
Learning control policies in simulation enables rapid, safe, and cost-effective development of advanced robotic capabilities. However, transferring these policies to the real world remains difficult due to the sim-to-real gap, where unmodeled dynamics and environmental disturbances can degrade policy performance. Existing approaches, such as domain randomization and Real2Sim2Real pipelines, can improve policy robustness, but either struggle under out-of-distribution conditions or require costly offline retraining. In this work, we approach these problems from a different perspective. Instead of relying on diverse training conditions before deployment, we focus on rapidly adapting the learned policy in the real world in an online fashion. To achieve this, we propose a novel online adaptive learning framework that unifies residual dynamics learning with real-time policy adaptation inside a differentiable simulation. Starting from a simple dynamics model, our framework continuously refines the model using real-world data to capture unmodeled effects and disturbances, such as payload changes and wind. The refined dynamics model is embedded in a differentiable simulation framework, enabling gradient backpropagation through the dynamics and thus rapid, sample-efficient policy updates beyond the reach of classical RL methods like PPO. All components of our system are designed for rapid adaptation, enabling the policy to adjust to unseen disturbances within 5 seconds of training. We validate the approach on agile quadrotor control under various disturbances in both simulation and the real world. Our framework reduces hovering error by up to 81% compared to -MPC and 55% compared to DATT, while also demonstrating robustness in vision-based control without explicit state estimation.
Supplementary Materials
Video: https://youtu.be/euK2GbcNTvk
Website: https://rpg.ifi.uzh.ch/lotf/
I Introduction
@@
Robot learning through simulation has seen great success in recent years, thanks to the rapid improvements in computer hardware and advancements in efficient physics simulation [collins2021review]. Simulation provides a fast, safe, and cost-effective way to collect data and train policies, enabling experiments that would be impractical or unsafe in the real world. However, transferring control policies learned purely in simulation to physical systems is challenging. While a high-fidelity simulation model may be used, the system parameters are often difficult to precisely identify. In addition, unmodeled effects such as aerodynamic turbulence, sensor noise, and actuator delays further complicate the real-world dynamics, thus making accurate alignment between simulation and reality difficult to achieve. The resulting mismatch, known as the sim-to-real gap [aljalbout2025reality], remains a central obstacle to deploying learning-based controllers in the real world. Bridging this gap is essential to retain the advantages of simulation while ensuring that policies perform reliably under real-world complexity and variability.
@@
Domain randomization is a common strategy to address this issue [DR], in which simulation parameters such as dynamics [kumar2021rma, kaufmann23champion], sensor noise [hwangbo2019learning, openai2019solving], and visual appearance [DR, xing2024contrastive] of the environment are varied during training to expose the policy to a wide range of possible deployment scenarios. By learning in diverse conditions, the agent develops robust policies that are less likely to overfit to the specific characteristics of a single environment. However, domain randomization cannot exhaustively anticipate all possible real-world conditions. Thus, when the environment shifts beyond the randomized distribution, policy performance will strongly degrade to out-of-distribution disturbances [wang2024environment]. Beyond domain randomization, Real2Sim2Real methods [hwangbo2019learning, kaufmann23champion] have shown strong sim-to-real transfer capabilities through offline refinement of the simulation model using real-world data before retraining policies for deployment. While such methods are effective in improving transfer, they require extensive data and costly retraining. For example, [bauersfeld2021neurobem] reports 75 minutes of real-world data collection, which makes them unsuitable for rapid online adaptation to changing conditions.
@@
In this work, we approach these problems from another perspective: we propose to rapidly adapt the policy to unknown external disturbances in the real world, in an online fashion. The core insight of the proposed framework is to integrate online residual dynamics learning with rapid policy adaptation via differentiable simulation. All system components are designed to update both the residual dynamics and policy as quickly as possible, ideally within a few seconds, during runtime. In this way, the policy becomes adaptive by continuously “overfitting” rapidly to the current environment scenario, and paradoxically, more “generalizable” across diverse conditions.
@@
For our pipeline, we start with a lightweight rigid-body dynamics model and continuously refine it by learning residual dynamics from real-world flight data. The residual-augmented dynamics model is embedded in a differentiable simulation framework to achieve more accurate and sample-efficient policy adaptation. Differentiable simulation provides the key advantage here: by allowing accurate gradients to flow through the dynamics, it makes real-world policy adaptation more efficient than classical RL approaches such as PPO [song2024learning]. Another key innovation is our alternating optimization scheme, where policy learning and residual model learning are interleaved so that each batch of real-world data is used efficiently for both dynamics refinement and control improvement through simulation using the refined dynamics. All of these components ensure the simulation is aligned with reality and enables rapid, efficient policy adaptation to unknown disturbances, even controlling directly from perceptual input.
@@
We evaluate our proposed framework in both simulation and real-world experiments across various environmental disturbance conditions on an agile quadrotor platform, whose nonlinear dynamics and sensitivity to aerodynamic effects make it a challenging benchmark for adaptive control [foehn2022agilicious]. In state-based control tasks such as hovering, where the policy receives as input full quadrotor state information, our method attains an average error of , an reduction over -MPC ( ) and over DATT [huang2023datt] ( ), while ensuring stable flight under modeling errors and out-of-distribution disturbances. In visual feature-based control, our framework achieves similar gains, demonstrating that rapid adaptation remains effective under partial or noisy observations - a capability unattainable with classical control methods in the absence of state estimation.
@@
This work aims to enable rapid, real-time policy adaptation on real robotic systems by integrating differentiable simulation and online residual dynamics learning. To the best of our knowledge, this is the first demonstration of coupling a differentiable simulator with real-time residual learning to achieve rapid, closed-loop policy adaptation in the real world. The framework continuously calibrates a lightweight analytical dynamics model with real-world data, allowing fast and stable adaptation to unseen disturbances. An alternating optimization scheme interleaves residual model updates and policy learning in a closed-loop manner, ensuring each real-world sample contributes to both model fidelity and control performance. Several design choices, including backpropagation through the analytical model only and an asynchronous implementation for concurrent data collection and optimization, were critical to achieving this speed and robustness. The framework supports both state-based and visual feature-based inputs (without explicit state estimation), and we demonstrate its effectiveness through simulated and real-world experiments, where it outperforms both classical and learning-based controllers under large unseen disturbances. Together, our framework demonstrates that policies can learn and adapt within seconds in the real world, thereby reducing reliance on domain randomization, which can fail to capture real-world complexity and adapt to out-of-distribution scenarios.
@@ II-A Aligning Simulation with the Real-World
Closing the sim-to-real gap requires quantifying the misalignment between simulation and real-world dynamics, typically through system identification or residual dynamics learning. System identification estimates parameters of an analytical dynamics model from input–output data [ljung1999system], but its representation capacity is limited to the modeled system [heiden2021neuralsim], making it less effective for capturing complex dynamics and disturbances. Residual dynamics learning addresses this by directly modeling the discrepancy between analytical predictions and real-world measurements. It has been applied to improve quadrotor odometry and tracking [bauersfeld2021neurobem, torrente2021data], learn motor delays in quadrupeds [hwangbo2019learning], and predict residual forces in soft robots [gao2024sim].
@@ II-B Fast Policy Learning in Simulation
While traditional RL methods suffer from prohibitively long training times, significant speed-ups can be achieved using highly optimized physics simulation [eschmann2024learning]. However, these methods are limited by sample inefficiency due to the high variance in the zeroth-order policy gradient estimates. A recent alternative paradigm, policy learning via differentiable simulation, uses smooth, differentiable dynamics and rewards to enable policy learning via first-order gradients [newbury2024review], offering substantial gains in sample efficiency and training time over RL [heeg2024learning]. It has been applied to direct policy parameterizations, such as parametric curve frequencies for swimming robots [nava2022fast] and sinusoidal policies for robotic cutting [heiden2021disect]. For applications to neural network policies, however, unstable gradients often limit applications to short-horizon tasks with simplified contacts and restricted start-state variation [xu2023efficient]. To address this, prior work has explored enhancements such as early-stopping simulations at contact events, truncated BPTT [xu2022accelerated], and reward augmentation with a learned critic [georgiev2024adaptive, luo2024residual].
@@ II-C Learning-Based Adaptive Control
Residual dynamics models have been used for online disturbance estimation via offline-trained networks [shi2019neural], Gaussian Processes [torrente2021data], or differentiable simulation–based system identification [chen2022real]. However, these methods mainly augment optimization-based controllers like MPC, which rely on full state information, and do not directly extend to vision-based control. Neural network policies have also been conditioned on disturbance estimates [huang2023datt, o2022neural, kumar2021rma], but since they are trained offline in randomized simulations and remain fixed during deployment, they struggle with domain shifts and unseen real-world conditions [xing2024multitask, huang2023datt].
@@ III Methodology
Our approach consists of two phases: policy pretraining and online adaptation. During pretraining, we train a base policy for online adaptation using a low-fidelity analytical dynamics model without residual dynamics. During online adaptation (see Fig. 3), residual dynamics learning, policy adaptation, and real-world deployment run in parallel across multiple threads, with parameters exchanged efficiently between processes via ROS as serialized byte strings. The residual dynamics network is continuously updated from a rolling buffer of quadrotor states and actions, and combined with the analytical model to form a hybrid dynamics embedded in the differentiable simulation for policy adaptation. The deployment loop uses the latest policy network parameters to continuously output control commands at 50 to the on-board flight controller, and simultaneously collects flight trajectories.
@@ III-A Differentiable Simulation Model
We model the quadrotor as a discrete-time dynamical system with continuous state and action spaces and , respectively. The system evolves according to the differentiable hybrid dynamics model which comprises the analytical and learned residual components, and describes the system evolution over time. At time step , an observation model generates an observation from the state , and is passed as input to a deterministic and differentiable policy network which outputs an action , and finally a deterministic, smooth and differentiable reward function emits a reward based on the state-action pair. Thus, all components are fully differentiable and allow gradient backpropagation through the simulation.
@@ III-C Policy Optimization Using Analytical Gradients
By leveraging the differentiable dynamics and reward structure, we can obtain first-order analytical policy gradients of the objective (2) via Back-Propagation Through Time (BPTT) (see [metz2021gradients] for a full derivation). The gradient and the update rule of the policy parameters are given by
where is the derivative matrix of the system dynamics , and is the learning rate. We build upon an existing open-source differentiable simulator for quadrotors [heeg2024learning] written entirely in JAX to leverage both its automatic-differentiation framework for computing the analytical policy gradients and performing GPU-accelerated parallel simulation.
@@ III-D Residual Dynamics Learning
Given the concatenated input vector of quadrotor state and action , an MLP network parameterized by is trained to predict the residual acceleration , defined as the difference between the ground-truth acceleration measured on the real system and the theoretical acceleration from the analytical dynamics in (1). The residual acceleration training targets are computed as . Given a batch of samples , we train the model by minimizing the loss function via where is the weight matrix of the -th network layer, and controls the regularization strength. The loss comprises a standard MSE term and a spectral norm regularization term, where the latter has been shown to improve generalization beyond the training distribution [bartlett2017spectrally] by regulating the network’s Lipschitz constant [shi2019neural].
@@ III-E Design Choices for Maximum Runtime Efficiency
During forward simulation, we use a hybrid dynamics model obtained by additively combining the analytical and learned residual dynamics models. Here, we use a simple, low-fidelity analytical dynamics model (see Sec. III-B) which models the quadrotor as a point-mass. The resulting acceleration given a state and action input pair is computed as , where is the network prediction. The quadrotor states are simulated at 50 via Runge-Kutta 4 time-integration of the dynamics using . While the hybrid dynamics model composed of differentiable analytical and learned components remains overall fully differentiable, we only perform gradient backpropagation through the analytical dynamics model and not the frozen network to obtain the policy gradients. This was inspired by prior work in policy learning using differentiable simulation for both quadruped [song2024learning] and quadrotor [heeg2024learning] control, which showed that combining accurate forward dynamics simulation with the backpropagation of a surrogate gradient based on a simplified dynamics model achieves faster runtime without impacting the resulting policy performance. We analyze and justify the above design choices through simulated experiments, and present and discuss the results in Sec. IV-C.
@@
We compare two existing methods of adapting a pretrained policy: full vs. low-rank adaptation (LoRA) [bo20243d]. Full adaptation involves updating all parameters of the policy network, similar to [xing2024bootstrapping], whereas LoRA freezes all pretrained parameters and instead adapts an additive low-rank network module [han2024parameter] which forms a lower-dimensional trainable parameter space. The latter has been shown to achieve more memory and parameter-efficient policy adaptation using RL for task-transfer [bo20243d] and multi-agent [zhang2025low] learning. Therefore, as an exploratory comparison, we seek to understand whether LoRA can also be combined with sample-efficient policy learning using differentiable simulation to effectively adapt a pretrained policy to unknown environmental disturbances.
@@ IV-A1 Task and Reward Definitions
We evaluate our approach on quadrotor stabilizing hover and trajectory tracking. For stabilizing hover, the policy is required to regulate the quadrotor towards a goal and maintain it at all times, which is non-trivial given the quadrotor’s non-linear and unstable dynamics. We evaluate both a state-based policy which receives observations at each time step, and an end-to-end visual feature-based policy which only receives the projected pixel coordinates of seven 3D keypoints from the past five time steps and the last three actions. For real-world experiments, the 3D keypoints are simulated in a hardware-in-the-loop style using quadrotor state estimates from a motion-capture system. Our training setup closely follows the open-source environment setup in [heeg2024learning]. Trajectory tracking requires following a reference trajectory defined as a time-parameterized sequence of quadrotor states, with policy training done in a state-based setting. As shown in Fig. 4, we generate two smooth trajectories, Circle and Figure-8, and a non-smooth 5-Point Star trajectory featuring a highly discontinuous velocity profile. For both tasks, the reward at each time step is defined as a sum of position, velocity, and actuation rewards For stabilizing hover, the individual reward terms are , and where is the Huber loss, and is the mass-normalized action required to counteract gravity. For trajectory tracking, the individual reward terms are and where and are respectively the reference position and velocity at time .